### v1

# MR-EEGWaveNet: Multiresolutional EEGWaveNet for Seizure Detection from Long EEG Recordings

Abstract
Feature engineering for generalized seizure detection models remains a significant challenge. Recently proposed models show variable performance depending on the training data and remain ineffective at accurately distinguishing artifacts from seizure data. In this study, we propose a novel end-to-end model, “Multiresolutional EEGWaveNet (MR-EEGWaveNet),” which efficiently distinguishes seizure events from background electroencephalogram (EEG) and artifacts/noise by capturing both temporal dependencies across different time frames and spatial relationships between channels. The model has three modules: convolution, feature extraction, and predictor. The convolution module extracts features through depth-wise and spatio-temporal convolution. The feature extraction module individually reduces the feature dimension extracted from EEG segments and their sub-segments. Subsequently, the extracted features are concatenated into a single vector for classification using a fully connected classifier called the predictor module. In addition, an anomaly score-based post-classification processing technique was introduced to reduce the false-positive rates of the model. Experimental results were reported and analyzed using different parameter settings and datasets (Siena (public) and Juntendo (private)). The proposed MR-EEGWaveNet significantly outperformed the conventional non-multiresolution approach, improving the F1 scores from $0.177$ to $0.336$ on Siena and $0.327$ to $0.488$ on Juntendo, with precision gains of $15.9\%$ and $20.62\%$, respectively.
Keywords: Epilepsy, Electroencephalogram (EEG), Seizure, Deep Learning, EEGWaveNet.
Acknowledgment: This work was supported in part by JSPS KAKENHI 23H00548. We would like to thank Editage (www.editage.jp) for English language editing.
1 Introduction
Epilepsy is a group of neurological disorders, characterized by unprovoked and recurrent seizures in a patient over a certain period [1]. A seizure is an outcome of abnormal activity in the brain. This abnormal activity can affect the entire brain or be confined to a particular brain region, determining the types of seizures [2]. Seizures can be caused by genetic disorders or brain injury; however, their underlying reasons often remain unknown [3]. The physiological and psychological effects of epilepsy in a patient can be reduced with an early and affordable diagnosis.
The electroencephalogram (EEG) is an essential tool for the diagnosis of epilepsy, as it captures electrical signals from the brain through electrodes [4]. In medical practice, scalp EEG is widely used for the primary inspection of seizure events. This process is non-invasive, affordable, and easily set up by practitioners. In addition, as seizures are rare events, their detection in EEG requires extensive data collection, often spanning hours, or days. Detecting seizures through an EEG is an important but cumbersome process for clinicians [5]. It requires higher-level expertise to distinguish seizure events from background EEG. Therefore, automatic seizure detection has been in high demand among medical professionals.
With advancement in technology and the availability of EEG data, numerous attempts have been made over the past few decades to automate the seizure detection system [5]. EEG data are inherently statistically variable across patients, making the feature engineering task quite challenging. Additionally, the nonlinear, non-stationary, and subject-dependent nature of EEG, including a heterogeneous recording environment, makes the task complex, especially in the case of scalp EEG recordings. Recent advancements in deep learning methods have inspired many researchers to develop models for detecting seizure events more efficiently [6, 7, 8, 9, 10, 11, 12]. Most existing models are based on convolutional neural networks (CNNs). However, several limitations exist in these studies, particularly regarding the use of hybrid architectures, validation techniques, and validation datasets. In hybrid architectures, feature extraction often partially depends on hand-crafted feature engineering. This approach may limit the generalizability of the model and fully utilize the deep learning representation capabilities. Furthermore, the K-fold cross-validation technique is a common validation approach to evaluate model performance. In epileptic datasets, considering the limited number of subjects and practical scenario, the leave-one-subject-out (LOSO) validation approach is more appropriate than K-fold cross-validation [13, 14]. In K-fold cross-validation, there is a high possibility that data from the same subject may appear in both the training and the testing sets, potentially leading to an overestimation of the model’s performance. Additionally, some publicly available EEG datasets [15], which are preprocessed to be noise-free and artifact-free, contain noninvasive and invasive EEGs together. When a model is validated on such datasets, it can significantly overestimate its true effectiveness in real-world scenarios. An end-to-end model called “EEGWaveNet” has demonstrated promising performance in the seizure detection task [16]. However, the performance of the model significantly depends on the training strategy and requires preprocessing to reduce the amount of noise or artifacts in the dataset. In contrast, the recording techniques for scalp EEG recordings are highly noisy or artifact-prone [17]. This can lead to frequent false seizure detections, undermining the model’s reliability.
In this study, we propose an extended version of “EEGWaveNet,” named “MR-EEGWaveNet”, which improves the performance of the original “EEGWaveNet” model by reducing the false detection of seizures for long EEG recordings. A key challenge for deep models in seizure detection lies in achieving balanced performance on highly imbalanced datasets, where accurate identification of the minority seizure class is critical. In contrast to the EEGWaveNet model, which uses fixed segment length, the MR-EEGWaveNet extracts features from the EEG segment and its sub-segments (multiple segments within the segment) and calculates their features for the classification task. The experimental results indicate that the MR-EEGWaveNet can obtain an optimal trade-off point of performance between recall and specificity, improving the precision score. We experimented with different parameter settings and performed a comparative analysis.
The major contributions of this paper are as follows:
  1. i.
    Proposed an extended version of the “EEGWaveNet” considering multiresolutional analysis in the feature extraction module.
  2. ii.
    A novel post-classification processing technique based on an anomaly score calculated by a state-of-the-art method further minimizes the misdiagnosis of the nonseizure EEG segment as seizure.
The performance of the MR-EEGWaveNet is evaluated with a public dataset (“Siena”) and a private dataset (“Juntendo”). While direct comparison with existing methods is difficult due to differing parameters, evaluation strategies, and datasets, we performed a comparative analysis with a recent state-of-the-art model. The results indicate that the MR-EEGWaveNet performs better in several key evaluation metrics.
The paper is organized as follows. Section 2 presents recent and relevant studies in the seizure detection task. Section 3 briefly discusses the experimental datasets. Section 4 contains the details of the proposed model. Sections 5, 6, and 7 describe the experimental setup, report the results, and discuss the findings, respectively. Finally, Section 8 provides concluding remarks.
2 Previous Works
Several studies have previously explored potential solutions to the seizure detection problem [18, 19, 20, 16]. They proposed several models and extracted features to distinguish seizure and nonseizure events from long recordings of EEG. In EEG research, the calculated features used to separate brain activity are commonly categorized into three groups: complexity, continuity, and connectivity [21]. In epilepsy research, complexity features (such as entropy and fractal-based methods) measure the irregularity in the data to detect seizure activity in EEG. Recently, different types of entropy have been proposed in the seizure detection task [22, 23]. Continuity-based features, such as band power ($\delta$, $\theta$, $\alpha$, $\beta$, and $\gamma$), median frequency, diffuse slowing, and burst suppression [24], calculate the stability and persistence of EEG signals over time. Connectivity features, such as coherence, phase-locking value (PLV), and mutual information, assess the relationship and interaction between EEG channels that cover different regions of the brain [21]. Furthermore, a combination of time-, frequency-, and entropy-based features is applied with a random forest classifier to detect seizures in EEG [25].
Although these features have been used in several studies, challenges and limitations, such as sensitivity to noise, longer time window requirements, limited temporal resolution, capture of linear relationships in EEG, and complex interpretation of brain connectivity, are encountered in the practical domain [21]. In addition, artifacts are common in scalp EEG recordings that reduce the performance of seizure detection models. The statistical properties of artifact and seizure events are similar, and their probability of simultaneous occurrence is relatively high [17]. To address this issue, the Riemannian manifold-based methods have recently gained popularity in the brain–computer interfacing domain. The Riemannian manifold is a nonlinear space that can be constructed from EEG [26, 27]. The Riemannian features are calculated through covariance matrices generated from EEG signals [28]. These features can be used effectively to detect anomalies in EEG and separate artifacts and seizures [29, 30].
However, the nonlinear and nonstationary nature of EEG presents a significant challenge in extracting features that consistently perform well over time. In general, EEG signals exhibit substantial inter-subject variability, and the features extracted from them are often highly subject-specific. In recent years, researchers have been actively developing data-driven approaches, such as end-to-end deep learning models to overcome feature engineering challenges in EEG [20, 16]. Convolutional neural networks (CNN) and bidirectional long short-term memory are well-known models that are used together to detect seizures, as proposed in [20]. Using a discrete wavelet transform (DWT), the EEG signals are decomposed into subbands, and the features are extracted using CNNs. Subsequently, the false detection rate (FDR) and sensitivity were improved by smoothing and collar techniques. By transforming EEG into plot images, a CNN-based model for seizure detection was designed in [31]. The model was improved with the addition of a patient-specific autoencoder (AE), which introduced a new label, “nonseizure-but-abnormal,” to the existing binary labels (“seizure” and “nonseizure”) [32]. Recently, graph-based analysis has also gained popularity in seizure detection. In [33], the short-time Fourier transform (STFT) of EEG signals is transformed into graph data and then classified with Google-Net CNN models. In [34], the EEG signal was first transformed into a spectrogram using the Stockwell transform. A Linear Graph Convolution Network (GCN) was used for feature selection, followed by DenseNet-based classification. “EEGWaveNet” is a deep learning model to process EEG signals introduced in [16]. It can capture both time-related and channel-related features that are suitable for classifying EEG signals compared to other models. However, implementing deep learning-based models in the medical domain (specifically seizure detection) has faced several challenges. Due to the complex and subject-dependent nature of seizures, these techniques have faced challenges, such as extreme data imbalance, complex interpretation, and the need for large amounts of training data to fine-tune hyperparameters.
In the case of epileptic seizure detection tasks, post-classification processing is a common approach to improve the performance of the system [5, 35, 36]. In [35], the authors adopted a post-processing approach based on multiple thresholds. They classified an event as a seizure if its probability exceeded the seizure threshold or if a background event was shorter than the minimum acceptable duration. In addition, seizures with a duration shorter than the minimum threshold were treated as background activity. In [36], seizures detected using an XGBoost-based method with intervals shorter than $2$ s were grouped and reclassified as nonseizure events if their duration was less than $15$ s.
Detecting seizures in long-term scalp EEG recordings is particularly challenging due to the frequent appearance of abnormal activities and artifacts. Therefore, special care and checking are required both during and after EEG recording. The selection of window length plays an important role in seizure detection; more specifically, in deep learning-based methods, where techniques, such as convolution are used [31]. Larger windows are suitable for extracting statistical properties, whereas a shorter windows provide more temporal information. Seizures are a complex and subject-dependent process. Certain seizures show consistent characteristics within and between consecutive time frames, while others have evolution, that is, their seizure characteristics change over time [37]. In addition, certain types of seizures have results, such as blinking in the eye, rapid movement in the eye, muscle contractions, stiffening, rhythmic jerking, lip smacking, hand movements, or chewing [38]. Those artifacts originate outside the brain and severely affect the pattern or shape of the EEG signal. Thus, the context of those artifacts is important to consider when designing a seizure detection model. However, different types of seizures and artifacts require varying window lengths to be effectively represented and distinguished by the detection model. Certain artifacts and seizures are better characterized with short windows for fine temporal details, while others require longer windows to extract broader patterns. For example, a high-amplitude spike artifact within a 10-second window may generate features that are indistinguishable from those of a seizure event if adjacent temporal dynamics are not considered.
A potential solution to the aforementioned problem is the multiresolutional feature analysis using deep learning-based models. Among recently proposed end-to-end models, the simplicity and potential of the EEGWaveNet model in seizure detection inspire us to further investigate it for possible performance improvement through architectural enhancements. In this study, we propose a new architecture based on the EEGWaveNet model, named “MR-EEGWaveNet,” designed to improve the performance of the existing model. The model consists of three main modules: convolution, feature extraction, and predictor. In contrast to EEGWaveNet, additional features are calculated at multiple resolutions from the sub-segments of the input EEG segment. This model explores the relationship among different parts of the EEG signal through multi-scale and spatio-temporal convolutions, including pooling operations. For each segment and its sub-segment, the feature extraction module extracted features from each “Spatio-temporal Convolution” module and concatenated them. Finally, the features from the segments and sub-segments are concatenated together to feed into the predictor module for the classification. This additional feature extraction from multiple resolutions enhances the ability to capture fine details of the short-time event while preserving their broader context. In addition, we introduced an anomaly score-based post-classification processing technique to improve the performance of the classification model.
(a) Nonseizure EEG
(b) Seizure EEG
Figure 1: $10$-seconds long (a) Nonseizure, (b) Seizure EEG segments from Juntendo Dataset.
3 Dataset
3.1 Public Dataset
A publicly available EEG dataset was employed. The dataset was recorded at the Unit of Neurology and Neurophysiology of the University of Siena [39]; hereafter, it will be referred to as “Siena”. The dataset comprised recordings of 14 patients, including nine males (ages 25–71) and five females (ages 20–58). The International 10-20 standard was used for EEG recordings with the $512$ Hz sampling frequency. The EEG recordings were downsampled to $500$ Hz for the experiment. The dataset was diagnosed and labeled (seizure events) by an expert clinician according to the International League Against Epilepsy criteria [40]. This study was conducted using EEG recordings from 19 commonly used channels: Fp1, Fp2, F7, F8, F3, F4, T3, T4, T5, T6, C3, C4, P3, P4, Fz, Cz, Pz, O1, and O2. Detailed information and statistics of the dataset are summarized in Table 1. The dataset is publicly available and can be accessed in [40].
Table 1: Description of the “Siena” EEG dataset. (Pt.: Patient ID; IAS: focal onset impaired awareness, WIAS: focal onset without impaired awareness, FBTC: focal to bilateral tonic-clonic; Loc.: Localization, T: temporal lobe, F: frontal lobe; Lat.: Lateralization)
Pt. Gender
Age
Seizure
types
Loc. Lat.
Record
Length(hr.)
Num. of
Seizures
Seizure
time (sec.)
Seizure
ratio (%)
PN00 Male 55 IAS T Right 3.20 5 325 2.821
PN01 Male 46 IAS T Left 13.48 2 128 0.264
PN03 Male 54 IAS T Right 24.22 2 244 0.280
PN05 Female 51 IAS T Left 6.02 3 104 0.480
PN06 Male 36 IAS T Left 12.05 5 282 0.650
PN07 Female 20 IAS T Left 8.73 1 62 0.197
PN09 Female 27 IAS T Left 6.83 3 203 0.825
PN10 Male 25 FBTC F Bilateral 18.68 10 365 0.543
PN11 Female 58 IAS T Right 2.40 1 55 0.637
PN12 Male 71 IAS T Left 6.07 4 290 1.328
PN13 Female 34 IAS T Left 8.63 3 264 0.849
PN14 Male 49 WIAS T Left 20.43 4 163 0.222
PN16 Female 41 IAS T Left 4.85 2 230 1.317
PN17 Male 42 IAS T Right 5.10 2 153 0.833
3.2 Private Dataset
In this study, a private EEG dataset recorded at Juntendo University Hospital in Japan was employed. The ethics committees of Juntendo University Hospital and Tokyo University of Agriculture and Technology approved the research. The dataset consists of scalp EEG recordings from multiple epileptic patients, with several hours of data collected for each. Specifically, 21 patients’ data (fourteen males and seven females), ranging in age from 1 to 42 years, were used in the experiment. The electrode placement followed the International 10-20 system, with a sampling frequency of 500 Hz. The number of selected channels was 19, the same as the Siena EEG dataset. The dataset contains different types of seizures, such as generalized, motor, nonmotor, focal, non-focal, tonic-clonic, behavioral arrest, among others. The brief data statistics are shown in Table 2. The $10$-s long nonseizure and seizure EEG segment from the dataset is shown in Figure 1.
Table 2: Description of the “Juntendo” EEG dataset (Pt.: Patient ID; Age y.: Year, m.:Month; hr.: Hours, sec.: Second)
Pt. Gender
Age
(y. m.)
Seizure
types
Channels with
seizure
Record
length (hr.)
Seizure
time (sec.)
Seizure
ratio (%)
1 Female 19y 10m
Generalized,
Motor, Tonic-clonic
All Channels
except Fp1,F7
4 217 1.507
2 Male 16y 5m
Generalized, Motor,
Epileptic spasms
C3, F3, T3 4 247 1.715
3 Male 26y 11m
Focal, Nonmotor,
Behavior arrest
F7, T3 8 371 1.288
4 Male 1y 4m
Generalized, Motor,
Epileptic spasms
All Channels 2 17 0.236
5 Male 8y 5m
Focal, Nonmotor,
Behavior arrest
C3, F3 6 441 2.042
6 Male 30y 9m
Focal, Nonmotor,
Behavior arrest
F7, T3 6 416 1.926
7 Male 27y 0m
Focal, Nonmotor,
Behavior arrest
F8, T4 4 161 1.118
8 Male 26y 4m
Focal, Motor,
Automatisms
F7, T3 4 229 1.59
9 Female 19y 4m
Generalized, Motor,
Tonic-clonic
All Channels
except Fp1, F7
2 1150 15.972
10 Male 24y 7m
Focal, Focal to
bilateral tonic-clonic
F7, T3 10 218 0.606
11 Female 27y 8m
Focal, Motor,
Automatisms
T5 4 343 2.382
12 Male 11y 7m
Focal, Motor, Tonic
F7,F8,
T3,T4,T5,T6
6.03 1581 7.282
13 Female 27y 9m
Focal, Motor,
Automatisms
All Channels
except Fz,Cz,Pz,
C1,C2
2 135 1.875
14 Male 35y 5m
Focal, Motor,
Automatisms
F7, F8 12 388 0.898
15 Female 5y 2m
Focal, Motor,
Epileptic spasms
C4, F4 4 125 0.868
16 Male 23y 10m
Generalized, Motor,
Tonic-clonic
All Channels 2 94 1.306
17 Male 42y 0m
Generalized, Motor,
Tonic-clonic
All Channels 2 23 0.319
18 Female 12y 11m
Unknown, Motor,
Epileptic spasms
C3, F3, F7, T3 4 59 0.41
19 Female 11y 8m
Generalized, Nonmotor,
Typical absence
All Channels 10 370 1.027
20 Female 35y 0m
Focal, Motor,
Automatisms
T3, T5 10 495 1.375
21 Male 31y 6m
Focal, Motor, Tonic
T3, T4, T5, T6 2 27 0.375
3.3 Preprocessing
As a preprocessing step, each EEG recording was filtered with an FIR bandpass filter from 1 to 60 Hz. In addition, a 50-Hz notch filter was applied to eliminate power noise from the recording system. The filtered EEG recordings were segmented with a fixed-length window (non-overlapping), with the window length depending on the experiment. In total, 19 commonly used monopolar channels were selected for the experiment (see Section 3.1).
4 Proposed Model
4.1 MR-EEGWaveNet: Multiresolutional EEGWaveNet
Table 3: MR-EEGWaveNet Architecture
Module Layer Kernel Output Activation
A Convolution Module
I) Multi-scale Convolution
Input ($C$, $N$) Linear
Conv1D kernel 2, stride 2, group $C$ ($C$, $N$/2) Linear
Conv1D kernel 2, stride 2, group $C$ ($C$, $N$/4) Linear
Conv1D kernel 2, stride 2, group $C$ ($C$, $N$/8) Linear
Conv1D kernel 2, stride 2, group $C$ ($C$, $N$/16) Linear
Conv1D kernel 2, stride 2, group $C$ ($C$, $N$/32) Linear
Conv1D kernel 2, stride 2, group $C$ ($C$, $N$/64) Linear
II) Spatio-temporal Convolution
Input ($C$, $N$/$2^{k}$)
Conv1D (32$\times$4) (32, $N$/$32^{k}-$ 3)
BatchNorm1D
Activation Leaky ReLU
Conv1D (32$\times$4) (32, $N$/$62^{k}-$ 6)
BatchNorm1D
Activation Leaky ReLU
Global average pooling 32
B Feature Extraction Module
Input 160
Fully Connected 64 Leaky ReLU
Fully Connected 32 Sigmoid
C Predictor Module
Input K
Fully Connected 64 Leaky ReLU
Fully Connected 32 Sigmoid
Classifier 2 Log Softmax
(a) Convolution Module
(b) MR-EEGWaveNet Architecture
Figure 2: Classification process of MR-EEGWaveNet model a) Convolution Module, and b) Overall MR-EEGWaveNet Architecture (Here, $10$ s, and $2$ s represent $10$ seconds, and $2$ seconds long multichannel EEG signal)
The MR-EEGWaveNet model contains three main modules: convolution, feature extraction, and predictor. The “convolution” module has a series of convolutional layers and pooling operations to process EEG signals, and is divided into two parts: a multi-scale convolution module and a spatio-temporal convolution module. The “multi-scale convolution” module has six temporal convolutional layers, each reducing the signal sequence length while maintaining channel information. The output of each temporal convolutional layer feeds to the next convolutional layer (except the last convolutional layer). The “spatio-temporal convolution” module performs pooling operations across multiple blocks to extract features from all channels. In each block, there are three operational components: convolutional layers, batch normalization, and LeakyReLU activation. For the EEGWaveNet architecture, we used two blocks sequentially to calculate features. Here, the inputs are the outputs of the temporal convolutional layers from “multi-scale convolution” (except the first temporal convolutional layer). Thus, five “spatio-temporal convolution” modules were used in the proposed model, each producing an output of $1\times 32$. Subsequently, they were concatenated to form a vector of dimension $1\times 160$.
Thereafter, the “feature extraction” module shrank the dimension of the feature from $160$ to $32$ with two fully connected layers with LeakyReLU activation. Thus, the final output of modules A and B for an EEG segment was $32$. The same process applies to the sub-segments. For example, if the proposed model has multiresolutional parameters $10$ s and $2$ s, then it will calculate one feature vector ($1\times 32$) for $10$ s and five feature vectors corresponding to five sub-segments of $2$ s (each has the same dimension of $1\times 32$). After concatenation, the final size of the feature vector will be $1\times 192$. Here, the list of multiresolutional parameters defines the architecture of the model. Figure 2 shows the architecture for the parameter list “$10$ s, $2$ s”. In this case, the model subdivided the $10$-s EEG signal into $2$ s segments; in total, six EEG signals (one $10$ s and five $2$ s multichannel EEG signals) were generated. Each signal was fed to modules A through B to calculate the features of each of them. Finally, the concatenated feature vector was fed into the predictor module for classification. The predictor module contains a fully connected classifier that maps the features to the output classes. The class probabilities were computed using the log-softmax function. The convolution module and the detailed architecture of the MR-EEGWaveNet for the above example are shown in Figures 2(a) and 2(b), respectively.
The main difference between EEGWaveNet and MR-EEGWaveNet architecture is that MR-EEGWaveNet extracts features both from the EEG segments and their sub-segments using modules A and B (the convolution and feature extraction module). Thus, the EEG segment was segmented into multiple levels based on the parameters. The architecture of MR-EEGWaveNet is shown in Table 3.
4.2 Post-Classification Processing
To further improve the performance of the proposed architecture, we applied an anomaly score-based post-classification processing technique to reduce the false detection rate of seizures. Therefore, we calculated the anomaly scores of the EEG segments, $a_{i}$ ($i$ is the segment index), with one of the state-of-the-art anomaly detection methods named ECOD: Unsupervised Outlier Detection Using Empirical Cumulative Distribution Functions [41]. Finally, the following rule was applied for deciding a segment as seizure or nonseizure based on the anomaly score of the EEG segments:
$\mu_{a}$ $=\frac{1}{M}\sum_{i=1}^{M}a_{i}$ (1)
$L_{i}$ =$=\begin{cases}1,&\text{seizure , if }a_{i}>\mu_{a}\\ 0,&\text{nonseizure}.\end{cases}$ (2)
where $\mu_{a}$ is the mean of $a_{i}$, $M$ is the total number of segments, and $L_{i}$ is the predicted label of a segment.
5 Experiment
The raw EEG data from each patient were preprocessed and segmented into equal length, $W$ (window length in seconds), using a sliding window approach with no overlap. Each segment of EEG has a dimension of $C\times N$, where $C$ is the number of channels, and $N$ is the number of time points of a single channel. Let $F_{s}$ be the sampling frequency, then $N=WF_{s}$.
For EEGWaveNet, the feature dimension is $F$=$32$, whereas in the case of MR-EEGWaveNet, it depends on the multiresolution parameter list $D$ (such as [$10$ s, $5$ s, $2$ s]). The number of parameters and the computational complexity of the models depend on $D$, $F$, and $W$. Here, the concatenated feature length $K$ depends on them and can be calculated as:
$K=\sum_{d\in D}F\left\lfloor\frac{W}{d}\right\rfloor$ (3)
[…]
Six different models were evaluated on the datasets. For EEGWaveNet we have experimented with EEGWaveNet-1 [2 s], EEGWaveNet-2 [5 s], and EEGWaveNet-3 [10 s] models corresponding to window lengths of $2$ s, $5$ s, and $10$ s, respectively. In the case of the proposed model, the three experimented models are listed as follows: MR-EEGWaveNet-1 [5 s, 2.5 s], MR-EEGWaveNet-2 [10 s, 2 s], and MR-EEGWaveNet-3 [10 s, 5 s, 2 s].
The performance of the models was evaluated with the metrics: precision (Pre.), recall (Rec.), specificity (Spe.), F1 score (F1), false positive rate (FPR), seizure detection ratio (Det. Ratio), and receiver operating characteristic–area under the curve (AUC).
$\mathrm{Precision}$ $=\frac{TP}{TP+FP}$ (4)
True Positive Rate $\text{True Positive Rate or}\;\mathrm{Recall}$ $=\frac{TP}{TP+FN}$ (5)
$\mathrm{Specificity}$ $=\frac{TN}{TN+FP}$ (6)
[…]
$\mathrm{FPR}$ $=\frac{FP}{FP+TN}$ (8)
Det. Ratio $=\frac{\text{Number of detected seizures}}{\text{Total number of seizures}}$ (9)
where $TP$, $TN$, $FP$, and $FN$ represent the total number of correctly detected positive, correctly detected negative, incorrectly detected positive, and incorrectly detected negative EEG segments, respectively. Moreover, the Receiver Operating Characteristic–Area Under the Curve (ROC-AUC) is a standard metric used to evaluate the ability of a binary classification model to differentiate between two classes. The ROC curve plots the True Positive Rate (TPR) (also known as recall) against the FPR at various classification thresholds. The Area Under the Curve (AUC) measures the entire area beneath the ROC curve and quantifies the overall performance of a model. The AUC is a single value that ranges from 0 to 1. An AUC of 1.0 indicates perfect classification, while an AUC of 0.5 does not reflect discriminative ability, the same as random guessing. Generally, a classifier with a higher AUC is considered to be more effective in binary classification tasks. Finally, the Det. Ratio represents the seizure detection ratio from the outcome of a model.
The models were evaluated with a LOSO validation approach, where one patient was designated for testing and the rest for training. In the case of preparing a dataset for training, we adopted the following two techniques to overcome the data imbalance problem:
  1. I.
    Seizure data: Oversampling; the seizure segments (within the expert’s notation of the seizure label) are 80%-overlapped between consecutive segments.
  2. II.
    Nonseizure data: Undersampling; the nonseizure segments are randomly selected from interictal EEG [42], the ratio between seizure and nonseizure is 1:2. This selection ratio is maintained for all recording sessions individually (that is, for each recording file).
For the test set, the recordings were segmented with a non-overlapping sliding window approach with a fixed value of $W$, where each segment was assigned to a label either 0 (nonseizure, negative class) or 1 (seizure, positive class) based on the ground truth label provided by an expert.
We used the following parameter values to train each model: batch size: $32$, number of training epochs: $300$, and learning rate: $10^{-3}$. The cross-entropy loss function was employed with balanced weights of $0.75$ and $1.5$ for the nonseizure and seizure classes, respectively. A warm start strategy with an early stop was used, where training continues until the validation loss stops improving for consecutive epochs (patience=$30$). For each iteration, the training and validation sets were randomly sampled, and the test set remained fixed across all runs. Each experiment was conducted three times to get a stable result. For each patient, performance metrics were averaged across the three runs, that is, the final score for a patient was computed as the mean of their results over the three iterations.
6 Results
6.1 Experimental Objectives
The objective of our experiments was to investigate the impact of multiresolution parameters on the performance of the classification model. We considered the following aspects when selecting parameters for the experiments:
  1. i)
    Effect of Window Length on Model Performance: The model behavior with different window lengths.
  2. ii)
    Performance of the Proposed Model: Achieving better performance using the proposed model.
  3. iii)
    Contribution of Post-Classification Processing: Contributions made by the post-classification processing method to the models.
6.2 Effect of Window Length on Model Performance
To investigate the aspect i), we considered the window lengths of $2$ s, $5$ s and $10$ s as parameters in the EEGWaveNet model, and their corresponding models were EEGWaveNet-1 [2 s], EEGWaveNet-2 [5 s], and EEGWaveNet-3 [10 s], respectively. In the Siena dataset (see Table 4), performance improves consistently with longer window lengths. Among all, EEGWaveNet-3 achieves the best results with the $10$-s window, indicating the contribution of the longer window size to the seizure detection task. A similar improvement in performance (except for specificity) with longer window lengths can be observed in the Juntendo dataset (see Table 5). In this case, the specificity decreases by approximately 3% and 2.6% when the window length increases from 2 to 5 s and from 2 to 10 s, respectively.
Furthermore, we investigated our MR-EEGWaveNet with two different window length sizes and three different multiresolutional parameters, which are as follows: MR-EEGWaveNet-1 [5 s, 2.5 s], MR-EEGWaveNet-2 [10 s, 2 s], and MR-EEGWaveNet-3 [10 s, 5 s, 2 s]. The window length of model MR-EEGWaveNet-1 was $5$ s, where both MR-EEGWaveNet-2 and MR-EEGWaveNet-3 were $10$ s long. For the Siena dataset (see Table 4), all metrics except recall improved with longer window lengths, whereas recall decreased by approximately 2% in the MR-EEGWaveNet-2 and MR-EEGWaveNet-3 models compared to MR-EEGWaveNet-1. For the Juntendo dataset (see Table 5), all improved metrics remain almost the same with longer window lengths.
Table 4: Performance comparison between EEGWaveNet and the MR-EEGWaveNet on the Siena dataset.
Models [Segments] Pre. Rec. Spe. F1 AUC
EEGWaveNet-1 [2 s] 5.82 77.92 83.63 0.098 0.8614
EEGWaveNet-2 [5 s] 8.05 84.84 84.25 0.126 0.9097
EEGWaveNet-3 [10 s] 11.72 85.99 86.04 0.161 0.9188
MR-EEGWaveNet-1 [5 s, 2.5 s] 15.58 84.02 90.43 0.218 0.9291
MR-EEGWaveNet-2 [10 s, 2 s] 26.74 82.39 93.86 0.319 0.9421
MR-EEGWaveNet-3 [10 s, 5 s, 2 s] 25.48 82.06 93.15 0.316 0.9441
Table 5: Performance comparison of EEGWaveNet and MR-EEGWaveNet on the Juntendo Dataset.
Models [Segments] Pre. Rec. Spe. F1 AUC
EEGWaveNet-1 [2 s] 18.20 79.82 85.93 0.253 0.8823
EEGWaveNet-2 [5 s] 18.51 82.04 83.03 0.254 0.8962
EEGWaveNet-3 [10 s] 19.07 88.83 83.35 0.274 0.9266
MR-EEGWaveNet-1 [5 s, 2.5 s] 36.13 76.91 91.99 0.428 0.9279
MR-EEGWaveNet-2 [10 s, 2 s] 42.92 77.94 91.60 0.474 0.9414
MR-EEGWaveNet-3 [10 s, 5 s, 2 s] 43.75 78.60 92.61 0.488 0.9410
6.3 Performance of the Proposed Model
A comparative analysis between EEGWaveNet and MR-EEGWaveNet is presented in Tables 4 and 5 for the Siena dataset and the Juntendo dataset, respectively. As shown in Table 4, the proposed MR-EEGWaveNet achieves better performance than the baseline EEGWaveNet on the Siena dataset. While EEGWaveNet improves slightly with longer segment lengths, MR-EEGWaveNet has considerably higher precision (up to 26.74%) and specificity (up to 93.86%). Although the recall of MR-EEGWaveNet is slightly lower than EEGWaveNet-3 (82.39% vs. 85.99%), it achieves higher F1 scores (up to 0.319) and AUC values (up to 0.9441). Similarly, as presented in Table 5, MR-EEGWaveNet shows better performance over EEGWaveNet on the Juntendo dataset. In this case, MR-EEGWaveNet outperforms EEGWaveNet with significantly higher precision (up to 43.75%), specificity (up to 92.61%), F1 scores (up to 0.488), and AUC values (up to 0.9414). However, MR-EEGWaveNet shows lower recall than EEGWaveNet but achieves a better balance with higher specificity (92.61%) compared to EEGWaveNet’s 83.35%, which is considerably lower.
Table 6: Performance comparison between EEGWaveNet and the MR-EEGWaveNet, with and without the “post-classification processing” technique, on the Siena dataset.
Models [Segments] Pre. Rec. Spe. F1 AUC
EEGWaveNet-1 [ 2 s] 5.82 77.92 83.63 0.098 0.8614
EEGWaveNet-1 [ 2 s] (post-process) 7.21 65.55 90.55 0.118
EEGWaveNet-2 [5 s] 8.05 84.84 84.25 0.126 0.9097
EEGWaveNet-2 [5 s] (post-process) 9.59 81.71 89.64 0.149
EEGWaveNet-3 [10 s] 11.72 85.99 86.04 0.161 0.9188
EEGWaveNet-3 [10 s] (post-process) 12.80 82.35 89.67 0.177
MR-EEGWaveNet-1 [5 s, 2.5 s] 15.58 84.02 90.43 0.218 0.9291
MR-EEGWaveNet-1 [5 s, 2.5 s] (post-process) 18.93 81.16 92.89 0.252
MR-EEGWaveNet-2 [10 s, 2 s] 26.74 82.39 93.86 0.319 0.9421
MR-EEGWaveNet-2 [10 s, 2 s] (post-process) 28.70 79.66 95.05 0.336
MR-EEGWaveNet-3 [10 s, 5 s, 2 s] 25.48 82.06 93.15 0.316 0.9441
MR-EEGWaveNet-3 [10 s, 5 s, 2 s] (post-process) 27.48 79.49 94.92 0.335
Table 7: Performance comparison between EEGWaveNet and the MR-EEGWaveNet, with and without the “post-classification processing” technique, on the Juntendo dataset.
Models [Segments] Pre. Rec. Spe. F1 AUC
EEGWaveNet-1 [ 2 s] 18.20 79.82 85.93 0.253 0.8823
EEGWaveNet-1 [ 2 s] (post-process) 23.69 78.68 92.47 0.315
EEGWaveNet-2 [5 s] 18.51 82.04 83.03 0.254 0.8962
EEGWaveNet-2 [5 s] (post-process) 22.72 80.86 91.19 0.302
EEGWaveNet-3 [10 s] 19.07 88.83 83.35 0.274 0.9266
EEGWaveNet-3 [10 s] (post-process) 23.56 87.91 90.94 0.327
MR-EEGWaveNet-1 [5 s, 2.5 s] 36.13 76.91 91.99 0.428 0.9279
MR-EEGWaveNet-1 [5 s, 2.5 s] (post-process) 39.46 76.64 94.76 0.455
MR-EEGWaveNet-2 [10 s, 2 s] 42.92 77.94 91.60 0.474 0.9414
MR-EEGWaveNet-2 [10 s, 2 s] (post-process) 44.18 77.57 94.51 0.488
MR-EEGWaveNet-3 [10 s, 5 s, 2 s] 43.75 78.60 92.61 0.488 0.9410
MR-EEGWaveNet-3 [10 s, 5 s, 2 s] (post-process) 46.07 78.26 95.40 0.509
6.4 Contribution of the Post-Classification Processing
[…]
6.5 Feature Visualization
(a) EEGWaveNet
(b) MR-EEGWaveNet
Figure 3: t-SNE visualization of features calculated from EEG recordings from the EEGWaveNet and the MR-EEGWaveNet. (Dataset: Juntendo, Patient ID: $12$)
Furthermore, we used the t-distributed stochastic neighbor embedding (t-SNE) technique to visualize the distribution of features in two dimensions, as shown in Figure 3. The features were extracted from the final layer of both the EEGWaveNet and the MR-EEGWaveNet.
7 Discussion
7.1 Effects of Window Length in Seizure Detection
The experimental results in Tables 4 and 5 indicate that the longer window tends to capture more characteristics of seizures than the shorter window. Moreover, it improves specificity in most cases, i.e., it efficiently separates noise, artifacts, and nonseizure EEG segments from seizure EEG. Based on our analysis and previous studies [22, 1], seizure EEG exhibits higher signal regularity compared to nonseizure EEG. With longer window, this regularity can be more effectively captured by the models. However, the longer window has less specific temporal information or features compared to a shorter window. Common artifacts, such as eye blink, eye movement, chewing, loose electrode, among others have some regularity, but their duration is significantly short, approximately $250$ ms to $3$ s [43]. In contrast, some artifacts, such as loss of all electrode contact, muscle movement, or body movement, have long-duration effects on the EEG [17]. The experimental results indicate that the proposed architecture, MR-EEGWaveNet, effectively mitigates this issue. In addition, the EEG monitoring guidelines suggest a display of $10$ to $20$ seconds/page for routine EEG checkup [44].
7.2 Effects of Post-Classification Processing
An essential aspect of post-classification processing is selecting an appropriate threshold. Anomaly or outlier detection methods generally assign higher anomaly scores to artifacts, noise, or seizure EEG segments compared to background EEG activity [29, 30]. During seizures, synchronous neuronal activity across the whole brain or specific regions often causes the data distribution to deviate significantly from the mean anomaly score. Unless the recording is heavily contaminated with physiological or extra-physiological artifacts, the average anomaly score during seizures typically deviates significantly from the mean baseline. Although certain parts of a seizure may resemble normal EEG, they may be parts of the evolution of the seizure. Based on this observation, the mean anomaly score was chosen as the threshold, with experimental results supporting its effectiveness. However, more studies with larger datasets are necessary to determine the optimal threshold for real-world implementation.
7.3 Performance Improvement with MR-EEGWaveNet
The experimental results on the Siena dataset (see Table 4) show that the MR-EEGWaveNet outperformed the EEGWaveNet model in most evaluation metrics, although it achieved a slightly lower recall. In particular, the MR-EEGWaveNet-2 and MR-EEGWaveNet-3 configurations achieved the highest performance, indicating a more robust detection capability. These results highlight the effectiveness of the MR-EEGWaveNet with multiresolution analysis compared to the existing EEGWaveNet model. Similar to the Siena dataset, the experimental results on the Juntendo dataset (see Table 5) also demonstrate the superiority of the MR-EEGWaveNet model over the EEGWaveNet model. However, the recall of the MR-EEGWaveNet is slightly lower than that of the EEGWaveNet, but in these cases, the specificity of the EEGWaveNet is considerably lower, approximately $83\%$.
The patient-wise performance of the EEGWaveNet-3 and MR-EEGWaveNet-2 on the Siena dataset with post-classification processing is shown in Tables A1 and A2, respectively. Furthermore, Tables B1 and B2 represent the results on the Juntendo dataset for the EEGWaveNet-3 and MR-EEGWaveNet-2 methods, respectively. Notably, the MR-EEGWaveNet demonstrates consistent performance across multiple datasets, further supporting its scalability and stability. However, more studies with larger datasets are necessary to confirm this finding.
Additionally, Figure 3 illustrates that the features calculated by the MR-EEGWaveNet (Figure 3(b)) exhibit better class-wise separation than those produced by the EEGWaveNet (Figure 3(a)), indicating an improved discriminative capability.
7.4 Selection of MR-EEGWaveNet Model
According to the performance of the proposed models in Table 6 and 7, both MR-EEGWaveNet-2 [10 s, 2 s], and MR-EEGWaveNet-3 [10 s, 5 s, 2 s] performs similar and better than MR-EEGWaveNet-1 [5 s, 2.5 s] in terms of precision, F1 score, and ROC-AUC. We conducted a Wilcoxon signed rank test using the Siena dataset to compare EEGWaveNet-3 with the MR-EEGWaveNet-2 and MR-EEGWaveNet-3 models, as all three models share the same window length. The p-value between EEGWaveNet-3 and MR-EEGWaveNet-2 is $0.0203$ $(<0.05)$, indicating a statistically significant performance difference, and MR-EEGWaveNet-2 outperforms EEGWaveNet-3 with a high degree of confidence. Furthermore, it indicates that the observed improvement is unlikely owing to random chance.
7.5 Challenges in Comparing Seizure Detection Systems
Performance comparison between seizure detection systems is difficult owing to differences in datasets, evaluation metrics, and experimental setups. The experimental datasets vary in terms of patient counts, seizure types, and recording settings, making the results difficult to compare. In addition, evaluation strategies and performance metrics differ across studies. Experimental protocols also vary, for example, the patient-dependent vs. independent approach. The preprocessing steps vary; some studies use artifact and noise removal techniques to clean the data for further processing. The definition of seizure events can also vary between datasets. Differences in expert opinion may arise when classifying an EEG segment as a seizure, particularly at the onset and offset of the event. Furthermore, the lack of explainability of the implementation of the methods reduces the reproducibility of the published work. While certain systems support offline analysis, real-time applications introduce added constraints, particularly regarding latency. Moreover, some studies consider segment-based evaluation, while others are event-based.
7.6 Comparison with Recent Studies
The Siena dataset utilized in this study is a relatively new dataset available publicly in the field. To the best of our knowledge, among the limited number of recent studies utilizing the Siena data set, the work by Sigsgaard et al. [25], which adopts a leave-one-subject-out evaluation protocol, represents the most comparable baseline for evaluating our approach. Although the method proposed in [25] is segment-based, the performance evaluation reported in their study is event-based. Table 8 presents a performance comparison between MR-EEGWaveNet-2 and the event-based approach proposed by Sigsgaard et al. In terms of performance based on seizures, our proposed method outperforms the approach of Sigsgaard et al. [25] by approximately $13\%$ in the detection of seizures, with a seizure event detection rate of $\sim 94\%$. However, in terms of specificity, it falls short by approximately $4\%$. Therefore, our proposed method effectively detects seizures in both segment-wise and event-based evaluation schemes, while maintaining a moderate level of specificity and a false positive rate. Furthermore, note that models designed explicitly for seizure event detection generally exhibit higher specificity compared to segment-based approaches. However, the performance reported largely depends on how seizure detection is defined and evaluated. In this study, a seizure event is considered detected if the model labels at least one EEG segment within the seizure period as a seizure, regardless of whether all segments of the event are labeled. For example, if an annotated seizure occurs between 100 and 150 seconds and the EEG is divided into 10-second segments, the event is considered detected if the model classifies any overlapping segment (e.g., 100–110 s or 110–120 s) as a seizure, even if not all segments within the 100–150 s period are detected.
Table 8: Performance comparison of the MR-EEGWaveNet with state-of-the-art works on the Siena Dataset.
Methods Recall Specificity
Sigsgaard et al. [25] 81.43 99.38
MR-EEGWaveNet-2 94.24 95.05
7.7 Performance Metric Trade-offs in Seizure Detection
[…]
7.8 Limitations and Future Directions
A notable limitation of our model is its reduced performance for certain subjects. Further improvements are needed to enhance the model’s ability to distinguish seizures from artifacts, such as loss of contact with the electrode, eye movements (EOG), or muscle activity. In the future, the model architecture should include context-based feature extraction to better capture temporal and spatial patterns. Furthermore, improving the model’s overall reliability and performance will require the application of advanced artifact removal and post-classification processing techniques.
8 Conclusion
This study presents a novel end-to-end model to detect epileptic seizures from long EEG recordings. The length of the EEG segment plays a significant role in the model’s performance, especially in distinguishing between seizure, noise, and artifact. The experimental results showed that the MR-EEGWaveNet with multiresolutional features performed better than the existing end-to-end models in classifying seizure and nonseizure EEG signals. Additionally, we introduced an anomaly score-based post-classification processing technique to further improve performance and reduce false positive rates. Furthermore, we demonstrated the performance of the MR-EEGWaveNet with different parameter settings. Despite lower performance in certain subjects, the improvement achieved with MR-EEGWaveNet indicates its potential for further development. However, enhancing MR-EEGWaveNet is essential to improve its ability to effectively distinguish between artifacts and seizure events. Although the MR-EEGWaveNet was tested on two datasets and demonstrated consistent performance, further analysis and investigation with larger datasets are required to apply it in real-world applications.
Appendix A Siena Dataset: Patient-Wise Results
A.1 EEGWaveNet Model
Table A1: EEGWaveNet-3, Patient-wise performance (Window length: $10$ s), with post-classification processing on Siena dataset (Pt.ID: Patient ID; Pre.: Precision; Rec.: Recall; Spe.: Specificity; F1: F1 score); The performance for each metric is reported as the average of three runs per patient.
Pt.ID Pre. Rec. Spe. F1 AUC Det. Ratio
Pt-00 45.25 100.00 96.95 0.622 0.996 1.000
Pt-01 3.92 100.00 87.62 0.074 0.970 1.000
Pt-03 1.71 86.07 87.47 0.033 0.881 1.000
Pt-05 1.42 58.52 81.66 0.028 0.903 0.889
Pt-06 18.55 80.77 96.52 0.285 0.967 0.867
Pt-07 0.95 80.00 84.82 0.019 0.907 1.000
Pt-09 44.73 83.33 95.57 0.406 0.975 0.889
Pt-10 17.04 35.32 98.40 0.201 0.683 0.407
Pt-11 5.24 94.44 87.74 0.098 0.914 1.000
Pt-12 3.74 91.22 70.81 0.072 0.901 1.000
Pt-13 10.74 92.06 93.52 0.191 0.958 1.000
Pt-14 2.78 69.05 95.19 0.053 0.911 0.750
Pt-16 6.71 89.39 83.08 0.124 0.917 1.000
Pt-17 16.43 92.68 95.99 0.276 0.942 1.000
Mean 12.80 82.35 89.67 0.177 0.918 0.914
A.2 MR-EEGWaveNet: Multiresolutional EEGWaveNet
Table A2: MR-EEGWaveNet-2, Patient-wise performance (Window: $10$ s), with post-classification processing on Siena dataset (Pt.ID: Patient ID; Pre.: Precision; Rec.: Recall; Spe.: Specificity; F1: F1 score); The performance for each metric is reported as the average of three runs per patient.
Pt.ID Pre. Rec. Spe. F1 AUC Det. Ratio
Pt-00 86.96 91.03 99.68 0.888 0.998 1.000
Pt-01 30.89 100.00 98.11 0.405 0.999 1.000
Pt-03 2.39 90.91 90.30 0.047 0.952 1.000
Pt-05 2.15 55.56 89.35 0.041 0.966 1.000
Pt-06 47.09 76.25 99.52 0.581 0.991 1.000
Pt-07 2.05 87.78 92.24 0.040 0.963 1.000
Pt-09 73.78 88.41 99.81 0.801 0.995 1.000
Pt-10 13.02 40.07 97.50 0.162 0.778 0.444
Pt-11 6.85 83.33 90.28 0.122 0.827 1.000
Pt-12 9.76 89.59 89.04 0.175 0.945 1.000
Pt-13 13.88 94.32 95.32 0.241 0.981 1.000
Pt-14 4.70 62.27 97.24 0.086 0.918 0.750
Pt-16 11.03 68.18 92.33 0.187 0.912 1.000
Pt-17 97.22 87.55 99.98 0.921 0.965 1.000
Mean 28.70 79.66 95.05 0.336 0.942 0.942
Appendix B Juntendo Dataset: Patient-wise Results
B.1 EEGWaveNet Model
Table B1: EEGWaveNet-3, Patient-wise performance (Window: $10$ s), with post-classification processing on Juntendo dataset (Pt.ID: Patient ID; Pre.: Precision; Rec.: Recall; Spe.: Specificity; F1: F1 score); The performance for each metric is reported as the average of three runs per patient.
Pt.ID Pre. Rec. Spe. F1 AUC Det. Ratio
Pt-01 50.27 93.65 98.10 0.633 0.995 0.922
Pt-02 4.35 100.0 66.32 0.083 0.909 1.000
Pt-03 20.62 99.05 94.30 0.335 0.989 1.000
Pt-04 5.22 100.0 91.55 0.098 0.997 1.000
Pt-05 6.81 83.47 81.45 0.126 0.798 1.000
Pt-06 37.42 98.29 96.13 0.529 0.995 1.000
Pt-07 31.21 97.92 96.25 0.454 0.992 1.000
Pt-08 15.02 95.38 90.90 0.258 0.956 1.000
Pt-09 74.07 83.55 94.17 0.779 0.887 1.000
Pt-10 29.51 74.20 98.97 0.417 0.911 1.000
Pt-11 25.51 80.81 94.27 0.384 0.960 1.000
Pt-12 46.65 66.11 93.81 0.529 0.811 1.000
Pt-13 5.99 85.71 72.19 0.112 0.906 1.000
Pt-14 23.58 86.20 96.39 0.352 0.989 1.000
Pt-15 4.62 97.44 81.88 0.088 0.826 1.000
Pt-16 49.92 96.30 98.45 0.644 0.998 1.000
Pt-17 2.88 66.67 90.44 0.053 0.895 0.667
Pt-18 2.91 76.67 87.89 0.056 0.870 1.000
Pt-19 22.44 96.40 95.34 0.356 0.974 0.957
Pt-20 9.30 68.23 91.64 0.163 0.814 1.000
Pt-21 26.45 100.00 99.21 0.417 0.995 1.000
Mean 23.56 87.91 90.94 0.327 0.927 0.978
B.2 MR-EEGWaveNet: Multiresolutional EEGWaveNet
Table B2: MR-EEGWaveNet-2, Patient-wise performance (Window: $10$ s), with post-classification processing on Juntendo dataset (Pt.ID: Patient ID; Pre.: Precision; Rec.: Recall; Spe.: Specificity; F1: F1 score); The performance for each metric is reported as the average of three runs per patient.
Pt.ID Pre. Rec. Spe. F1 AUC Det. Ratio
Pt-01 71.99 83.57 99.46 0.765 0.993 0.804
Pt-02 4.31 92.28 68.67 0.082 0.886 1.000
Pt-03 44.98 94.20 98.42 0.601 0.991 1.000
Pt-04 29.52 100.0 99.26 0.450 0.999 1.000
Pt-05 43.78 55.81 98.75 0.473 0.919 0.556
Pt-06 51.22 90.77 98.24 0.647 0.983 1.000
Pt-07 74.21 97.92 99.63 0.844 0.998 1.000
Pt-08 18.91 88.89 93.89 0.312 0.958 1.000
Pt-09 95.63 79.45 99.28 0.866 0.970 0.889
Pt-10 67.17 44.01 99.90 0.532 0.986 0.867
Pt-11 49.08 78.12 97.99 0.593 0.981 1.000
Pt-12 50.13 51.43 95.83 0.487 0.877 1.000
Pt-13 4.65 76.83 71.00 0.088 0.835 1.000
Pt-14 34.47 70.35 98.32 0.423 0.986 1.000
Pt-15 5.40 94.44 85.15 0.102 0.816 1.000
Pt-16 96.67 100.00 99.95 0.982 0.996 1.000
Pt-17 2.44 50.00 91.45 0.046 0.780 0.500
Pt-18 2.84 46.67 94.22 0.054 0.902 0.500
Pt-19 76.00 74.81 99.63 0.724 0.989 0.731
Pt-20 15.38 59.34 95.83 0.240 0.920 1.000
Pt-21 88.89 100.00 99.95 0.933 0.998 1.000
Mean 44.18 77.57 94.51 0.488 0.941 0.897